Pound Industries is attempting to select the best of three mutually exclusive projects. The initial investment and​ after-tax cash inflows associated with these projects are shown in the following table. Cash flows Project A Project B Project C Initial investment​ (CF) ​$60000 ​$100000 ​$110000 Cash inflows​ (CF), t equals1 to 5: ​$20000 ​$31500 ​$32500

a.  Calculate the payback period for each project.

b.  Calculate the net present value​ (NPV) of each​ project, assuming that the firm has a cost of capital equal to 13​%.

c.  Calculate the internal rate of return​ (IRR) for each project.

d.  Indicate which project you would recommend.
Answer #1

a. Initial Investment of A =60000
Cash flow from year 1 to year 5 =20000
Payback period formula = Years before recovery + Cost not covered in that year/ Cash flow for that year =60000/20000 +0 =3 years

Initial Investment of B =100000
Cash flow from year 1 to year 5 =31500
Payback period formula = Years before recovery + Cost not covered in that year/ Cash flow for that year
=3 +(100000-31500*3)/31500=3.17 years

Initial Investment of C =110000
Cash flow from year 1 to year 5 =32500
Payback period formula = Years before recovery + Cost not covered in that year/ Cash flow for that year
=3 +(110000-32500*3)/32500=3.38 years

b. NPV of Project of A =PV of Cash flow -Initial investment =20000*(((1-(1+13%)^-5)/13%)-60000=10344.63
NPV of Project of B =PV of Cash flow -Initial investment =31500*(((1-(1+13%)^-5)/13%)-100000=10792.78
NPV of Project of C =PV of Cash flow -Initial investment =32500*(((1-(1+13%)^-5)/13%)-110000=4310.02

c. IRR of A using Financial Calculator
CF0=-60000;CF1 =20000;CF2 =20000;CF3 =20000;CF4 =20000;CF5  =20000;CPT IRR =19.86%

IRR of B using Financial Calculator
CF0=-100000;CF1 =31500;CF2 =31500;CF3 =31500;CF4 =31500;CF5  =31500;CPT IRR =17.34%

IRR of C using Financial Calculator
CF0=-110000;CF1 =32500;CF2 =32500;CF3 =32500;CF4 =32500;CF5  =32500;CPT IRR =14.59%

d. Project B should be accepted because it provides highest NPV.
​​​​​​​